OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95.

OpenStudy (anonymous):

Tn= a1 +(n-1)d for nth term there a18-a1= a1+(18-1)d from here you can calculate d and then you will have n=35 and d put it in Tn and you will have your answer

OpenStudy (anonymous):

@lostandconfused1

OpenStudy (anonymous):

203?

OpenStudy (anonymous):

@theyatin

OpenStudy (anonymous):

you have d=6 and n=35

OpenStudy (anonymous):

then Tn=-7+(35-1)6 =-7+34*6=-7+204=197